# AI-Enhanced Self-Triggering for Extensive Air Showers: Performance and FPGA Feasibility

### v1

@@ Abstract
Cosmic-ray detection with radio antennas has traditionally depended on external triggers from particle detectors, constraining sensitivity and increasing complexity. Previous attempts at fully standalone, radio-only triggers have often failed under intense radio frequency interference, making genuine air-shower signals difficult to isolate. We present a proof-of-principle artificial intelligence-based self-triggering system that overcomes these limitations. By training a deep learning model on both real noise data and injected cosmic-ray-like pulses, we achieve an exceptionally low false-positive rate alongside high detection efficiency. Configurable operating points can suppress false positives below 0.01% while retaining more than 88% of genuine signals, and can even eliminate false positives entirely at a modest reduction in signal efficiency. This flexibility makes single-station cosmic-ray detection feasible without requiring external trigger inputs. Applying our approach to real-world noise conditions reduces the initial false-positive event rate by several orders of magnitude, supporting large-scale deployments. Extrapolation to dedicated hardware implementations, such as FPGAs, indicates that sub- inference times are achievable, enabling real-time autonomous triggering. These results highlight the transformative potential of artificial intelligence for enhancing radio detection sensitivity and inaugurate a new generation of fully self-triggered cosmic-ray observatories.
@@ 1 Introduction
Recent developments in modelling EAS radio emissions, combined with experimental and simulation studies, have provided a deeper understanding of characteristic pulse shapes and ground radiation patterns [2]. This progress has paved the way for more sophisticated signal-processing techniques, including artificial intelligence (AI)-driven approaches for enhancing radio-based self-triggering. Recent studies show that deep learning models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), can effectively distinguish cosmic-ray signals from background noise, significantly reducing detection thresholds. These techniques have been successfully applied to identify air-shower radio pulses in offline analyses [10], demonstrating their potential to enhance signal selection. Recent efforts, such as those in the GRAND experiment, continue to investigate self-triggering using AI-based methods, including CNNs. While preliminary results indicate that these methods can reject at least 40% of background noise at a 90% signal selection efficiency [11], achieving real-time deployment remains challenging. Self-triggering requires ultra-low latency processing to ensure the identification of cosmic-ray signals within microseconds of their arrival, while also maintaining a false-positive trigger rate at the sub- level, a constraint that traditional software-based neural networks struggle to meet.
@@
In this work, we investigate the feasibility of FPGA-based AI self-triggering for radio cosmic-ray detection. We begin by training and evaluating our deep neural network on Metal Performance Shaders (MPS) with data that includes experimentally measured background noise, then estimate their potential performance on FPGAs. While high-level synthesis (HLS) frameworks such as HLS4ML [14] and Xilinx Vitis AI [15] offer streamlined approaches to deploying AI models on FPGAs, we consider their integration a logical extension of this proof-of-concept study rather than part of our immediate scope.
Our method comprises a fully convolutional autoencoder that denoises radio traces by suppressing background noise while retaining cosmic-ray pulses, followed by a fully convolutional classifier (FCN) that differentiates cosmic-ray signals from noise. By leveraging real data with measured background noise, we enhance the robustness of our approach in practical deployments. We also address critical issues associated with FPGA-based AI, including quantization strategies, resource utilization, and latency reduction, and propose a framework for real-time classification of cosmic-ray radio pulses. This proof-of-principle demonstration paves the way for future work incorporating more advanced HLS techniques to further optimize deployment on FPGA hardware.
The implementation of AI-accelerated FPGA-based self-triggering represents a significant advancement in radio detection technology. By reducing reliance on external detector triggers and increasing detection efficiency, this approach enables greater autonomy for radio observatories, facilitating the detection of rare and high-energy events, including ultra-high-energy neutrinos. Our results demonstrate the feasibility of deploying AI-driven self-triggering at the detector level, marking an important step toward fully autonomous, real-time radio detection systems for astroparticle physics.
@@ 2 Experimental Setup
The dataset used in this study was collected in an urban environment, where background noise levels are higher and more variable compared to remote cosmic-ray observatories. To capture a diverse range of real-world noise conditions, data was collected over approximately two hours, yielding 200 million samples. Data was continuously streamed, with traces read out at a rate of one per second. Given the high variability of the urban noise environment, the dataset presents a challenging test case for AI-driven signal detection methods. This setting introduces additional challenges, including increased transient interference and strong radio frequency interference (RFI). Two prominent narrowband RFI sources were identified at approximately and , along with other narrowband and transient RFI. The power spectrum of the recorded traces, shown in Figure 1, was obtained after applying a digital bandpass filter in the range of 30 to . For the initial analysis, these narrowband RFIs were not suppressed to assess the model’s robustness in the presence of strong interference. The results presented in the following section reflect model performance when both training and testing were conducted on data sets containing these RFIs. Additionally, we examine how performance changes when these narrowband RFIs are mitigated prior to training and validation, providing insights into the impact of RFI suppression on the learning process, which we will report in the discussion section 7.
To evaluate the AI model’s performance, signal-like pulses were injected into measured noise traces at random positions, covering a broad range of amplitudes. The pulse power was scaled relative to the noise power, which was calculated using a pre-filtered version of each trace. This ensured that the injected pulses maintained a well-controlled signal-to-noise ratio (SNR) within a user-defined range while accurately reflecting real-world noise conditions. However, to prevent double filtering of the noise, which could distort its characteristics, the pulses were injected into the original (unfiltered) traces. This step also eliminated systematic biases by ensuring that the statistical properties of the background noise remained unaffected by the injection process. To maintain consistency in data processing, the digital bandpass filter was applied only after pulse injection, ensuring that both noise-only and signal-injected datasets underwent identical filtering and conditioning. This approach was chosen over COREAS simulations [16] to establish a proof of principle based purely on signal strength, avoiding dependencies on the absolute energy scale, antenna gain pattern, and shower inclination, which introduce additional uncertainties. By injecting short pulses into measured noise traces from a real-world, high-interference environment, we enable a focused evaluation of the AI model’s ability to detect transient signal shapes under conditions where traditional threshold-based triggering systems struggle or fail entirely due to overwhelming radio frequency interference (RFI). This approach enables a systematic assessment of the model sensitivity in varying SNRs while ensuring robustness under challenging conditions. By directly testing and refining AI-based techniques on measured noise traces—without relying on an external event-reconstruction framework or requiring a particle detector—it provides a controlled and scalable test case. Although it does not fully replicate all details of cosmic-ray pulse shapes, much of which may already be washed out by bandpass filtering, it offers a robust estimate of the false positive rate of AI-based triggering methods in a realistic noise environment.
@@ 2.1 Training and Validation Datasets
To train and evaluate the model, multiple datasets were used, each designed to fulfil a specific role in ensuring robust performance. The datasets were categorized into training, cross-validation, and independent validation sets to assess the model’s generalization and sensitivity. All noise traces are unique and are derived from experimental measurements, randomly selected from the full duration of the noise measurement campaign, whereas the signal-containing traces consist of injected pulses, as previously described. For the initial analysis, only a bandpass filter was applied, with no additional filtering. However, we later incorporate further filtering to suppress prominent RFI sources in the 30– range for comparison, the results of which are discussed in section 7.
@@
Training Dataset (50%): Comprised 15,000 time series traces, each containing 128 samples recorded at a sampling rate of , consistent with standard practices in radio-based air shower detection. It included an equal distribution of 5,000 pure noise traces (background), 5,000 traces of isolated EAS-like pulses (pure signal), and 5,000 traces with EAS-like pulses injected into noise (signal). This setup ensured that the model effectively learned to distinguish genuine signals from background noise.
@@
Figure 2 illustrates the SNR distributions for background and signal traces across Validation Dataset 1, which closely follows the distribution of the training dataset, and Validation Dataset 2, which consists of low-amplitude signals. This comparison provides insights into the differences in signal detectability across varying SNR conditions.
@@ 3 Methodology
We implemented a deep learning-based pipeline for denoising and classifying radio time series in the context of self-triggering for EAS events. The pipeline consists of two primary components:
1.
A CNN-based denoiser to suppress noise in the input signals.
2.
A fully convolutional classifier (FCN) to distinguish between signal and background in a binary classification task.
These models operate sequentially, where the denoiser first refines raw data before passing it to the classifier for event discrimination.
@@ 3.1 Model Architecture
3.1.1 CNN-Based Denoiser
The denoiser is a fully convolutional autoencoder designed to suppress noise and enhance cosmic ray pulses in noisy radio time series.. It follows an encoder-bottleneck-decoder structure with a skip connection to retain fine-grained features.
Encoder: Two convolutional layers extract multi-scale features, followed by a single max pooling layer that reduces resolution by a factor of two.
Bottleneck: A deeper convolutional block refines feature representations at the reduced resolution.
Decoder: An upsampling layer restores the original resolution, followed by convolutional layers that refine the denoised signal. A skip connection concatenates high-resolution features from the encoder to enhance detail recovery.
The denoiser learns from a dataset where the inputs are noisy traces containing EAS-like pulses (signal), and the targets are the corresponding pure EAS-like pulses (pure signal).
The training process minimises the Mean Squared Error (MSE) loss between the clean target and the denoised prediction :
This architecture ensures that the model preserves cosmic ray features while effectively filtering out background noise.
3.1.2 Fully Convolutional Classifier (FCN)
The FCN classifier classifies time series into background or signal, where signal refers to a noisy trace containing an EAS pulse. It includes:
Three convolutional blocks with batch normalization, dropout, and max pooling.
Global average pooling (GAP) to reduce spatial dimensions while retaining features.
A fully connected layer with a sigmoid activation function for binary classification.
The classifier is trained using the Binary Cross-Entropy (BCE) loss:
where is the true label and is the predicted probability.
3.2 Training and Optimization
Both models were trained using the Adam optimizer, with task-specific loss functions, following a two-step training approach:
1.
Denoiser Training:
Trained on traces with pulses injected into noise (signal) and pure signal.
Optimised to minimize Mean Squared Error (MSE) loss, enhancing signal quality while suppressing noise.
2.
Classifier Training:
Trained on denoised signal traces (output of the denoiser) alongside denoised background traces.
Optimized to minimize Binary Cross-Entropy (BCE) loss, ensuring accurate discrimination between signal-containing and noise-only traces.
The model was trained for 15 epochs with a batch size of 64 and a learning rate of 0.001, using the Training Dataset. Its generalization performance was monitored with the Cross-Validation Dataset described in Section 2.1. The model was implemented using PyTorch with support for MPS acceleration. Training was conducted using the Adam optimizer, and model performance was evaluated on the cross-validation dataset using classification accuracy, precision, recall, and F1-score.
4 Model Performance Analysis
The first two plots in Figure 3 (top row) illustrate the probability distributions of sigmoid prediction for validation datasets 1 and 2, which differ in their SNR distributions. In Validation Dataset 1, where the SNR distribution closely resembles that of the training set (see Section 2.1), the model exhibits a strong separation between background and signal traces. Most noise traces receive lower probabilities, while signal-containing traces are predominantly assigned higher values. In contrast, Validation Dataset 2 contains traces with significantly lower SNRs, commonly below the noise floor. While the model still distinguishes between the two classes, the classification performance degrades: the distribution of signal-containing traces is more spread out, with a noticeable tail extending toward lower probabilities. This shift indicates a reduction in signal efficiency, as weaker signals are increasingly misclassified due to their low SNR.
Beyond assessing performance across different SNRs, it is also essential to evaluate the model’s behaviour in environments dominated by overwhelming background noise, where false positives become a critical concern. To investigate this, we introduce Validation Dataset 3, which consists of one million background traces. The bottom row of Figure 3 explores the model’s response to this dataset by comparing background traces from Validation Dataset 3 with signal traces from Validation Dataset 1. The sigmoid probability distribution (bottom left) shows that while most noise traces receive low probabilities, the large dataset size results in a fraction extending into the signal region. This highlights the need to quantify the false trigger rate under realistic high-background conditions. The final plot (bottom right) presents the distributions of the model’s logits, i.e. the raw scores before applying a probability function. These distributions reveal a clear separation between background traces from Validation Dataset 3 and signal traces from Validation Dataset 1. Using logits instead of sigmoid probabilities significantly improves false positive rejection, albeit at the cost of true positive efficiency. This trade-off is particularly relevant when designing self-triggering strategies in noisy environments.
4.1 Performance Metrics and Trigger Rates
Following the qualitative assessment in Figure 3, we systematically evaluate the model’s performance across different validation datasets. This analysis focuses on three key aspects: trigger efficiency, robustness to low-SNR signals, and false trigger rate in high-background conditions. The results are presented as follows.
4.1.1 Trigger Efficiency in Controlled Conditions (Validation Dataset 1)
True Positives: 99.93%
False Positives: 0.07%
A simple probability threshold is sufficient for efficient signal selection in this case.
4.1.2 Robustness to Weaker Pulses (Validation Dataset 2)
True Positives: 87.96%
False Positives: 0.08%
Despite the increased challenge, the model effectively distinguishes signal-containing traces, demonstrating its robustness to weaker pulses. The reduction in efficiency is expected as most of these signals fall outside the training range. Although further tuning could enhance performance, the primary objective here is to assess the model’s ability to generalise rather than optimise for specific low-SNR cases.
4.1.3 False Trigger Rate Estimation in a High-Noise Environment (Validation Dataset 3)
To assess the model’s performance in the presence of overwhelming background noise, we evaluated it on Validation Dataset 3. This dataset simulates realistic high-RFI conditions, such as those found in urban environments (e.g., LOFAR, LOPES) or certain AERA stations [7].
The primary objective of this evaluation is to quantify false trigger rates in such environments. The results are as follows:
True Positives (Initial Threshold): 99.93%
False Positives: 768 per 1M events (0.08%), which translates to approximately 800 Hz first-level trigger (FLT) rate in a 1 MHz trigger scenario.
Applying confidence-based selection criteria (referred here as entropy cut) significantly reduces the false positive rate while maintaining high true positive efficiency.
Entropy Cut: 99.35% True Positives, 79 False Positives per 1M events (0.008%), corresponding to approximately 80 Hz false triggers in a 1 MHz trigger scenario.
The FLT rate is already suitable for a station-level trigger and provides a strong foundation for further refinement. A subsequent second-level trigger (SLT) stage, utilizing array-level coincidence searches, could further suppress false positives while preserving efficiency.
To explore the feasibility of detecting cosmic-ray events at a single station, we further applied logits-based cuts, which resulted in:
Logit Cut: 72.4% True Positives, 0 False Positives, completely eliminating false triggers in this dataset.
This analysis shows that entropy-based selection effectively reduces the trigger rate to an acceptable level for an FLT, while still maintaining high detection efficiency. Further refinement at the SLT stage, through array-level coincidence searches, could optimise false positive rejection even further. On the other hand, applying logits-based cuts enables the possibility of detecting cosmic-ray events at a single station, drastically suppressing false positives, with none observed in our dataset. While this comes with some reduction in signal efficiency, it remains within a reasonable range, making single-station detection feasible in certain scenarios.
4.1.4 Example of a Low-Amplitude Signal Classification

### v3

@@ Abstract
Autonomous self-triggering for radio detection of extensive air showers remains a long-standing challenge, particularly in environments dominated by strong and variable radio-frequency interference. Current radio arrays usually rely on external particle-detector triggers: while this lowers thresholds for vertical showers, it excludes very inclined events, where radio detection is uniquely powerful because the particle cascade is absorbed while the radio pulse remains measurable. In this work, I present a proof-of-principle study showing that deep-learning-based triggering can overcome these limitations, operating robustly under realistic high-interference conditions and within the strict latency constraints of large-scale observatories. Using measured noise traces combined with simulated cosmic-ray pulses, a fully convolutional network was trained and optimised for MHz-scale trigger-trial rates, and its performance was compared to a simple threshold-based trigger, which proved markedly less efficient at the same false-positive rate. The trained model was then quantised with HLS4ML and synthesised with Vitis HLS for multiple FPGA targets. The quantised implementation preserves the floating-point model’s physics performance (, ) while achieving microsecond-scale inference latencies. These results show that modern FPGA-based AI inference can provide low-latency, radio-only triggering for next-generation cosmic-ray observatories, particularly in regions with strong and variable radio backgrounds. This capability unlocks access to weak and highly inclined air-shower signals—even in noisy environments—thereby broadening the energy range, extending sky coverage, and opening the path toward ultra-high-energy neutrino detection with ground-based radio arrays.
@@ 1 Introduction
Notably, the ANITA [10] and TREND [11, 12] experiments successfully implemented autonomous radio self-triggering. ANITA, a balloon-borne radio interferometer, operated in the stratosphere, where the continuous background was comparatively quiet but impulsive anthropogenic RFI events were still present and were identified and rejected using polarisation information in subsequent analysis. TREND, by contrast, deployed a ground-based array in a remote, radio-quiet region of China, where ambient noise levels were exceptionally low. In both cases, comparatively favourable noise conditions were crucial in enabling effective self-triggering with simple threshold-based approaches—conditions that are significantly more challenging in the high-interference environments considered in this work.
Recent developments in modelling EAS radio emissions, combined with experimental and simulation studies, have provided a deeper understanding of characteristic pulse shapes and ground radiation patterns [2]. This progress has paved the way for more sophisticated signal-processing techniques, including artificial intelligence (AI)-driven approaches for enhancing radio-based self-triggering. Recent studies show that deep learning models, including convolutional neural networks (CNNs) and recurrent neural networks (RNNs), can effectively distinguish cosmic-ray signals from background noise, significantly reducing detection thresholds. These techniques have been successfully applied to identify air-shower radio pulses in offline analyses [13, 14, 15], demonstrating their potential to enhance signal selection. Recent efforts, such as those in the GRAND experiment, continue to investigate self-triggering using AI-based methods, including CNNs. While preliminary results indicate that these methods can reject at least 40% of background noise at a 90% signal selection efficiency [16], achieving real-time deployment remains challenging. Self-triggering requires ultra-low latency processing to ensure the identification of cosmic-ray signals within microseconds of their arrival, while also maintaining a false-positive trigger rate at the sub- level, a constraint that traditional software-based neural networks struggle to meet.
@@
In this work, I investigate the feasibility of FPGA-based AI self-triggering for radio detection of EAS. I begin by training and evaluating a fully convolutional network on GPU using a dataset that combines experimentally measured background noise with simulated cosmic-ray pulses generated by the MGMR3D framework [19, 20, 21, 22]. The trained model is then quantised with hls4ml [23, 24, 25] to fixed-point precision and synthesised using Vitis HLS 2023.2 [26] for multiple FPGA targets, allowing direct assessment of post-synthesis latency and resource utilisation.
By leveraging measured noise from a high-interference environment, the approach is designed for robustness in realistic deployment conditions. I address key considerations for FPGA-based AI triggers, including quantisation, latency optimisation, and DSP-constrained architectures, and demonstrate that sub-microsecond inference latencies are achievable within realistic resource budgets.
The implementation of AI-accelerated FPGA-based self-triggering represents a significant advancement in radio detection technology. By reducing reliance on external detector triggers and enabling high-efficiency, low-latency signal identification, this work opens the path toward autonomous operation of next-generation radio observatories. It marks the first demonstration, in the context of radio-detector self-triggering, of a complete workflow from model training to FPGA-synthesised deployment using hls4ml, providing a solid foundation for future developments in real-time cosmic-ray radio detection.
@@ 2 Experimental Setup
The dataset used in this study was collected in an urban environment, where background noise levels are higher and more variable compared to remote cosmic-ray observatories. To capture a diverse range of real-world noise conditions, data was collected over approximately two hours, yielding 200 million samples. Data was continuously streamed, with traces read out at a rate of one per second. Given the high variability of the urban noise environment, the dataset presents a challenging test case for AI-driven signal detection methods.
This setting introduces additional challenges, including increased transient interference and strong RFI. Two prominent narrowband RFI sources were identified at approximately and , along with other narrowband and transient RFI. The power spectrum of the recorded traces, shown in Figure 1, was obtained after applying a digital bandpass filter in the range of 30 to . For the initial analysis, these narrowband RFIs were not suppressed to assess the model’s robustness in the presence of strong interference. The results presented in the following section reflect model performance when both training and testing were conducted on data sets containing these RFIs.
To evaluate the AI model’s performance under realistic and diverse signal conditions, I generated air-shower radio pulses using the MGMR3D simulation framework [19, 20, 21, 22], which reproduces the time-domain waveform and frequency content expected from extensive air showers for a wide range of shower geometries and primary energies. The simulated dataset spans zenith angles from to in steps, uniformly distributed random azimuths, and primary energies between and . For each geometry, was varied randomly between 600 and to cover the natural spread observed in air-shower development. Signals were simulated for antenna positions at distances of to from the shower core, ensuring realistic variation in pulse amplitude, shape, polarity, and spectral content across the dataset. In total, this resulted in the generation of over 78,000 simulated radio pulses.
These pulses, already encompassing a wide range of shapes and amplitudes from the simulated variation in energy, geometry, values and antenna–core distance, were injected at random times into measured noise traces from a high-interference environment. To control the signal-to-noise ratio (SNR) within a user-defined range—without performing an absolute amplitude conversion from simulated electric-field strength to the measured detector response—I rescaled each simulated pulse relative to the noise power, calculated from a pre-filtered version of the corresponding trace. To preserve the statistical properties of the noise and avoid any distortion from repeated filtering, pulses were injected into the original (unfiltered) traces. The same digital bandpass filter was then applied to both noise-only and signal-injected traces, ensuring identical processing and conditioning across the dataset.
By combining realistic MGMR3D-generated air-shower pulses with measured noise, this approach enables a robust and representative evaluation of AI-based triggering performance. It captures the natural diversity of pulse shapes expected in real events while providing controlled SNR conditions. This ensures that the results are directly relevant for estimating the achievable false-positive rate under realistic and challenging RFI conditions.
@@ 2.1 Training and Validation Datasets
To train and evaluate the model, multiple datasets were used, each designed to fulfil a specific role in ensuring robust performance. The datasets were categorized into training, cross-validation, and independent validation sets to assess the model’s generalization and sensitivity. All noise traces are unique and are derived from experimental measurements, randomly selected from the full duration of the noise measurement campaign, whereas the signal-containing traces consist of injected pulses, as previously described. For the initial analysis, only a bandpass filter was applied, with no additional filtering.
@@
Training Dataset (50%): Comprised 15,000 time series traces, each containing 128 samples recorded at a sampling rate of , consistent with standard practices in radio-based air shower detection. It included an equal distribution of 5,000 pure noise traces (background), 5,000 traces of simulated EAS pulses (pure signal), and 5,000 traces with simulated EAS pulses injected into noise (signal). This setup ensured that the model effectively learned to distinguish genuine signals from background noise.
@@
To quantify the SNR in a way that is robust to real detector conditions, I adopt a trace-wise definition using pure signal and pure noise traces (background) as inputs. Specifically, I define SNR as the squared peak amplitude of the Hilbert envelope of a pure signal trace, divided by the variance of a matched pure noise trace. This avoids ambiguity over whether the "signal" contains background or not. Formally, I compute
where is the maximum amplitude of the analytic signal (obtained via the Hilbert transform) and is the noise power computed as the variance of a pure noise trace. This approach emphasises the prominence of the signal over the background without being sensitive to specific waveform shapes.
To provide a baseline for background-only traces, I apply the same method to pure noise traces, treating them as if they contained a signal. Specifically, I compute the Hilbert envelope of each noise trace and take the maximum value as a hypothetical signal peak, then divide by the noise variance. This procedure captures spurious fluctuations in the noise that might mimic signal-like features. As a result, I obtain a distribution of apparent SNR values for background, which are typically lower than true signals but may overlap in some cases. This overlap justifies the use of logarithmic scaling and motivates the application of probabilistic, rather than deterministic, classification thresholds.
Figure 2 illustrates the SNR distributions for background and signal traces across Validation Dataset 1, which closely follows the distribution of the training dataset, and Validation Dataset 2, which consists of low-amplitude signals. This comparison provides insights into the differences in signal detectability across varying SNR conditions. In the SNR histograms, I report the base-10 logarithm of the computed SNR.
@@ 3 Methodology
The goal of my approach is to design, train, and evaluate a neural network classifier capable of robustly detecting air-shower radio pulses in realistic, high-interference noise conditions, while ensuring compatibility with efficient FPGA deployment. The methodology comprises two main components: (1) the design of a fully convolutional network architecture optimised for both classification performance and hardware constraints, and (2) the training and optimisation procedure used to achieve high detection efficiency with minimal false-positive rate.
@@ 3.1 Model Architecture
The classification network is a fully convolutional 1-D architecture for radio traces with one input channel. It contains approximately trainable parameters and was designed to maintain classification performance comparable to earlier floating-point models while ensuring that each convolutional or linear layer contains fewer than weights—a practical constraint arising from Vitis HLS synthesis failures when unrolling larger layers due to internal 12-bit address limitations in the memory architecture of Vitis HLS, which restrict access to weight arrays exceeding 4096 entries.
This constraint facilitates efficient FPGA implementation in later stages of the work without excessive on-chip memory use or pipeline latency.
The network consists of three convolutional blocks followed by global average pooling and a fully connected output layer. Batch normalisation (BN) is applied after each convolution, ReLU activations follow all but the final convolution in each block, and dropout is applied after each pooling stage.
The first block applies a 1D convolution with kernel size 3, 1 input channel, and 16 output channels (),
where denotes the ReLU activation. This is followed by max-pooling with factor 2 and dropout with rate 0.3.
The second block begins with a 1D convolution of kernel size 3 that increases the number of channels from 16 to 32, followed by batch normalisation and ReLU activation:
This is followed by a residual sub-network composed of lightweight 1D convolutions:
where consists of two 1D convolutional layers with reduced intermediate channel width, each followed by batch normalisation and activation. This design reduces the number of parameters while preserving the benefits of residual learning. The block ends with max-pooling (factor 2) and dropout with rate 0.3.
The third block increases the number of channels from 32 to 64 using a bottleneck-style expansion module composed of three 1D convolutions:
where and are 1D convolutions with kernel size 1, and is a 1D convolution with kernel size 3. Let denote the batch normalisation applied after the convolution. The first two convolutions are each followed by batch normalisation and ReLU activation; the final convolution is followed only by batch normalisation. This design enables parameter-efficient widening by reducing the number of channels in the intermediate layers, avoiding the cost of a wide convolution. The block concludes with max-pooling (factor 2) and dropout with rate 0.3.
The classification head applies global average pooling across the temporal dimension, reducing the output tensor to a 64-dimensional feature vector by averaging across the temporal axis. A fully connected layer (641) produces the binary classification logit without an activation function, as the model is trained using a loss function that incorporates the sigmoid non-linearity.
This description provides all hyperparameters, activation placements, and custom-module definitions required to reproduce the model exactly.
3.2 Training and Optimisation
The previously described architecture was trained to classify traces as either pure noise (background) or containing simulated cosmic-ray pulses (signal), as defined in Section 2.1. The loss function was the binary cross-entropy with logits,
where is the model’s raw output (logit), the binary target label, the sigmoid function, and the mini-batch size.
The model was optimised using the Adam algorithm [27] as implemented in PyTorch, with a fixed learning rate of , and default values for the exponential decay rates and numerical stability parameter: , , and . Training was performed for 15 epochs, and the model state corresponding to the lowest validation loss was retained for evaluation.
The dataset used for model training and evaluation consisted of normalised, fixed-length radio traces drawn from two equally sized classes, corresponding to the training and cross-validation datasets defined in Section 2.1. The dataset was randomly shuffled and split into training and validation subsets. Training mini-batches were drawn from both classes in equal proportion to avoid class imbalance.
To improve robustness for hardware-friendly deployment, unstructured magnitude pruning was applied to all convolutional layers. A polynomial schedule gradually increased weight sparsity from zero to 50% between early and mid-training epochs:
where denotes the sparsity at epoch , is the target sparsity, and and are the pruning schedule boundaries. During training, pruning masks were dynamically updated, and after training completed, masks were removed to yield the final pruned model.
After each epoch, the network was evaluated on the validation set without gradient updates. The primary selection criterion for saving the best model was the validation loss. For monitoring purposes during development, the area under the receiver operating characteristic (ROC) curve (AUC) was also computed using a fixed classification threshold of 0.5.
The complete training pipeline, including dataset preparation, augmentation through pulse injection, pruning schedule, and validation, was implemented in PyTorch and executed on CUDA-enabled GPUs.
4 Model Performance Analysis (Float Model)
To evaluate the classifier’s final performance in a realistic triggering context, I consider signal traces from Validation Dataset 1 and background traces from Validation Dataset 3. Validation Dataset 1 contains noise traces with injected MGMR3D-simulated air-shower pulses at controlled SNRs matching the training distribution. Validation Dataset 3 consists of one million measured background traces recorded in a high-interference environment. The large sample size enables the evaluation of the model’s false-positive rejection efficiency in a highly challenging regime, where even a small false-positive rate would correspond to rates approaching the MHz range in a real-time self-trigger application.
The main performance indicator is the receiver operating characteristic (ROC) curve (Figure 3), obtained from classifier outputs for signal traces from Validation 1 and background traces from Validation 3. The solid curve shows the trade-off between the true positive rate (signal efficiency) and the false positive rate as the classification threshold is varied, with an area under the curve (AUC) of 0.969. The vertical red dashed line marks the chosen operating point at a false positive rate of , illustrating that the model maintains high efficiency while suppressing false triggers to extremely low rates.
While the ROC analysis focuses on the nominal SNR distribution, model performance was also evaluated on Validation Dataset 2, where most injected pulses have SNRs well below the noise floor. At a fixed false positive rate of , the signal efficiency decreases from on Validation 1 to on Validation 2. This drop reflects the expected trend toward lower classification scores for weaker signals. These results provide a first indication that the network retains some discriminative power in the low-SNR regime.
4.1 False Trigger Rate Estimation in a High-Noise Environment
To assess the model’s robustness in realistic high-noise conditions, I evaluated it on Validation Dataset 3. This dataset reflects high-RFI environments such as those encountered in urban settings (e.g., LOFAR, LOPES) or at certain AERA stations [7]. Such conditions are representative of operational scenarios in which a self-triggering station must continuously process high-rate background waveforms, potentially at sampling rates of up to trigger trials. In this regime, even a small false positive rate (FPR) translates into a substantial trigger load: for example, an corresponds to first-level trigger (FLT) rate per station.
From the ROC analysis (Figure 3), the model achieves a signal efficiency of at this target operating point. The FLT rate is already suitable for a station-level trigger in a cosmic-ray radio array and provides a strong foundation for further refinement. In practice, a subsequent second-level trigger (SLT) stage—based on array-level coincidence requirements—can further suppress false positives while preserving nearly all genuine events. Such multi-stage triggering strategies are widely used in cosmic-ray experiments (e.g., Auger, AERA, GRAND) and are essential for handling high raw data rates in a scalable and efficient manner.
For applications where further reduction of the FLT rate is desired, the threshold can be tightened to , corresponding to roughly at the station level. This operating point still retains a signal efficiency of and may be useful in scenarios where exceptionally low background rates are advantageous. In practice, already offers a good balance between efficiency and manageable trigger rates, with the lower-FPR setting available as an option when background suppression takes priority.
4.2 Example of a Low-Amplitude Signal Classification